OpenStudy (anonymous):

Given the equation −4 sqrt x-3 = 12, solve for x and identify if it is an extraneous solution. x = 0, solution is not extraneous x = 0, solution is extraneous x = 12, solution is not extraneous x = 12, solution is extraneous

OpenStudy (anonymous):

@kewlgeek555 @AllTehMaffs can you help at all?

OpenStudy (kewlgeek555):

I don't unerstand the equation. Is it:\[\sqrt[-4]{x ^{-3}}=12\]

OpenStudy (anonymous):

I think it'd be \[-4\sqrt{x-3}=12\]

OpenStudy (anonymous):

You get a complex solution

OpenStudy (anonymous):

That doesn't give the right solutions though. Hmm....

OpenStudy (anonymous):

if it's \[\sqrt[-4]{x^{-3}} = 12\] \[ (x^{-3})^{-1/4}=12\] \[(x^{-3})^{-1} = 12^4\] \[x^3=12^4\] that gives a weird answer :P I can't figure out where they're getting either 0 or 12, however :P